# Mountain Pass Solutions for an entire semipositone problem involving the Grushin Subelliptic Operator

### v1

@@ 1 Introduction
Let $f$ satisfy (f1)-(f2) and let $g\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ be positive. Then there exists $a_{1}>0$ such that for each $a\in(0,a_{1})$ (SP) admits a weak solution $u_{a}\in C(\mathbb{R}^{N})\cap D_{\gamma}(\mathbb{R}^{N})$ such that $u_{a}\geq 0$ a.e. in $\mathbb{R}^{N}$. If, in addition,
@@
    $d(z)^{N_{\gamma}-2}\int_{\mathbb{R}^{N}}\frac{g(\xi)}{d(z-\xi)^{N_{\gamma}-2}}\mathrm{d}\xi<\infty$, where $z\in\mathbb{R}^{N}\setminus\{0\}$,
then there exists $\tilde{a}\in(0,a_{1})$ such that $u_{a}>0$ a.e. in $\mathbb{R}^{N}$ for every $a\in(0,\tilde{a})$.
@@ 3 A functional setting for the Grushin operator
is the Sobolev critical exponent associated to $N_{\gamma}=m+(1+\gamma)\ell$, the homogeneous dimension associated to the decomposition $N=m+\ell$. Since the Sobolev-Gagliardo-Niremberg inequality proved in [25, Lemma 5.1]
@@
Let $\{u_{j}\}$ be a bounded sequence in $D_{\gamma}(\mathbb{R}^{N})$. We can suppose that $u_{j}\rightharpoonup 0$. For each $R>0$, we have the continuous embedding $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow\mathscr{H}_{\gamma}(B_{R}(0))$ [^3] For an open bounded set $\Omega\subset\mathbb{R}^{N}$, we denote by $\mathscr{H}_{\gamma}(\Omega)$ the completion of $C^{1}_{0}(\Omega)$ with respect to the norm $\lVert u\rVert_{\gamma}=\left(\int_{\Omega}\lvert\nabla_{\gamma}u\rvert^{2}\,{\mathrm{d}z}\right)^{1/2}.$ Since the embedding $\mathscr{H}_{\gamma}(B_{R}(0))\hookrightarrow L^{q}(B_{R}(0))$ is compact, see [23, Proposition 3.2], it follows that $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow L^{q}(B_{R}(0))$ is a compact embedding as well. Hence,
@@
for some subsequence. As $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow L^{2^{*}_{\gamma}}(\mathbb{R}^{N})$ is a continuous embedding, we have that $\{\lvert u_{j}\rvert^{q}\}$ is a bounded sequence in $L^{2^{*}_{\gamma}/q}(\mathbb{R}^{N})$. By the Brezis-Lieb Lemma, [8, Theorem 1], up to a subsequence if necessary
@@ 4 Existence of solutions for (SP)
Now we are in a position to apply the Mountain Pass Theorem [2, Theorem 2.1], which ensures the existence and uniform boundedness of solutions to the (SP) problem.
@@
Consider $a_{1}$, $\delta$ and $v$ as given in Lemma 4.1. For $a\in(0,a_{1})$, using Lemma 4.1 and 4.2, the hypotheses of the mountain pass theorem [2, Theorem 2.1] are verified. So there exists a non-trivial critical point $u_{a}\in D_{\gamma}(\mathbb{R}^{N})$ of $I_{a}$ satisfying
@@ 5 Properties of the solutions
Our next result establishes that the solution $u_{a}$ belongs to $L^{\infty}(\mathbb{R}^{N})$ if the parameter $a$ is sufficiently small. Using the iterative argument of [28, Lemma B.3] we prove that the solution $u_{a}\in L^{q}(\mathbb{R}^{N})$, for every $q\in[2^{*}_{\gamma},\infty)$, provided $a$ is small enough. Moreover the argument in [26, Theorem 3.1] can be applied to establish the boundedness of $u_{a}$.
@@
Step 3. The continuity of the solutions $u_{j}$ comes from the non-homogeneous Harnack inequality for $X$-elliptic operators established in [18, Theorem 5.5] combined with some standard arguments, see for example [19, Corollary 4.18] for further details. ∎
@@ 6 Positive solutions
Let $\Sigma=\{0\}\times\mathbb{R}^{\ell}$ be the degenerate set for the Grushin operator. If $\Omega$ is any bounded domain contained in $\mathbb{R}^{N}\setminus\Sigma$, then $\Delta_{\gamma}$ is a uniformly elliptic operator on $\Omega$. Hence $u\in C^{2}(\Omega)$ and either $u\equiv 0$ on $\Omega$ or $u>0$ in $\Omega$. In the former case, it follows from [23, Proposition 2.5] that $u\equiv 0$ on the connected component of $\mathbb{R}^{N}\setminus\Sigma$ which contains $\Omega$. Therefore $u\equiv 0$ on $\mathbb{R}^{N}\setminus\Sigma$.
Now, it is clear that $\Sigma$ has empty interior, and it follows from [21, Theorem 9, Chapter 1] that $\mathbb{R}^{N}\setminus\Sigma$ is dense in $\mathbb{R}^{N}$. Recalling that $u$ is a continuous function, we conclude that $u\equiv 0$ on $\Sigma$. To this stage the function $u$ might be zero in (the interior of) some connected component of $\mathbb{R}^{N}\setminus\Sigma$ and strictly positive in (the interior of) some other component. However this is again impossible, since [23, Proposition 2.5] applies again on any ball which meets two such components across $\Sigma$. Hence either $u\equiv 0$ on $\mathbb{R}^{N}$, or $u>0$ on $\mathbb{R}^{N}$. The former is not admissible due to the Lemma 5.3. ∎
@@
When $\gamma$ is an even integer, a very elegant proof of the positivity of $u$ can be devised. We follow an idea of [4]. When $\gamma$ is an integer, it is known that $-\Delta_{\gamma}$ satisfies the Hörmander condition. Recalling that $g>0$ and $f_{0}\geq 0$, we see that $\int_{\mathbb{R}^{N}}\nabla_{\gamma}u\nabla_{\gamma}\varphi\,{\mathrm{d}z}\geq 0$ for every $\varphi\in C_{0}^{\infty}(\mathbb{R}^{N})$ such that $\varphi\geq 0$. Hence $\int_{\mathbb{R}^{N}}u(-\Delta_{\gamma}\varphi)\,{\mathrm{d}z}\geq 0$ for every such $\varphi$. Recalling that $u$ is a continuous function, it follows that $u$ is a viscosity subsolution of the operator $-\Delta_{\gamma}$. The strong maximum principle proved in [3, Corollary 1.4] yields that either $u\equiv 0$ or $u>0$ in $\mathbb{R}^{N}$. The first option, again, is not admissible due to the Lemma 5.3.
Now we are going to show that $u_{j}\to u$ in $L^{\infty}(\mathbb{R}^{N})$ norm. We proceed using the same argument in [1, 5] and the fundamental solution for the Grushin operator.
@@
is a fundamental solution for the Grushin operator, see [14, Appendix B].
@@
$C^{-1}=(N_{\gamma}-2)\int_{\{z\mid d(z)=1\}}\frac{\lvert x\rvert^{2\gamma}}{\left(\lvert x\rvert^{2\left(1+2\gamma\right)}+\left(1+\gamma\right)^{2}\lvert y\rvert^{2}\right)^{1/2}}\,\mathrm{d}S,$
@@
Let us consider a spherical shell $\Omega=B_{R_{2}}\setminus\overline{B_{R_{1}}}$, with $0\leq R_{1}<R_{2}\leq+\infty$. If $u$ is a radial function, by [13, Eq. (2.4)] we get the formula
@@
$\lVert u_{j}-u\rVert_{\infty}\to 0$ as $j\to\infty$.
@@
By Proposition (6.1)
$u(z)=C\int_{\mathbb{R}^{N}}\frac{g(\xi)f_{0}(u(\xi))}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi\quad\hbox{a.e. in}\leavevmode\nobreak\ \mathbb{R}^{N}.$ (33)
@@
$u_{j}(z)=C\int_{\mathbb{R}^{N}}\frac{g(\xi)f_{j}(u_{j}(\xi))}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi\quad\hbox{a.e. in}\leavevmode\nobreak\ \mathbb{R}^{N}.$ (34)
Using (33) and (34) we estimate $\lvert u-u_{j}\rvert$ as follows:
@@
Therefore, (35) yields $u_{j}\to u$ in $L^{\infty}(\mathbb{R}^{N})$ as $j\to\infty$. ∎
@@
$u_{j}\geq 0$ a.e. in $\mathbb{R}^{N}$ for $j$ large enough.
@@
If (f3) and (g1) are satisfied then $u_{j}>0$ a.e. in $\mathbb{R}^{N}$.
@@
for some $M>0$. For $z\in\mathbb{R}^{N}\setminus\{0\}$, using (33) and (34), we get
@@
for any $R>0$. For any $R>0$ there exists $z\in\mathbb{R}^{N}$ such that $d(z)>2R+1$. Hence
@@
Therefore, by (37) there exists $j_{2}\in\mathbb{N}$ and $R>>1$ such that for $j\geq j_{2},\,u_{j}>0$ a.e. on $\mathbb{R}^{N}\setminus B_{R}$. Moreover, since $u$ is continuous, there exists $\eta>0$ such that $u>\eta$ on $\overline{B_{R}}$. Therefore, from Proposition 6.5, there exists $j_{3}\in\mathbb{N}$ such that for $j\geq j_{3},\,u_{j}>0$ a.e. on $\overline{B_{R}}$. Thus, by choosing $j_{4}=\max\left\{j_{2},j_{3}\right\}$, we see that for $j\geq j_{4},\,u_{j}>0$ a.e. on $\mathbb{R}^{N}$. This completes the proof. ∎

### v2

@@ 1 Introduction
Let $f$ satisfy (f1)-(f2) and let $g\in L^{1}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ be positive. Then there exists $a_{1}>0$ such that for each $a\in(0,a_{1})$ (SP) admits a weak solution $u_{a}\in C(\mathbb{R}^{N})\cap D_{\gamma}(\mathbb{R}^{N})$ such that $u_{a}\geq 0$ a.e. in $\{0\}\times\mathbb{R}^{\ell}$. If, in addition,
@@
    for all $z\in(\{0\}\times\mathbb{R}^{\ell})\setminus\{(0,0)\}$,
    $d(z)^{N_{\gamma}-2}\int_{\mathbb{R}^{N}}\frac{g(\xi)}{d(z-\xi)^{N_{\gamma}-2}}\mathrm{d}\xi<\infty,$
then there exists $\tilde{a}\in(0,a_{1})$ such that $u_{a}>0$ a.e. in $\{0\}\times\mathbb{R}^{\ell}$ for every $a\in(0,\tilde{a})$.
@@
To the best of our knowledge, there are no results in the literature about semipositone problems for degenerate equations of Baouendi-Grushin type. Before concluding this introduction, we would like to comment about the positivity in the subspace $\{0\}\times\mathbb{R}^{\ell}$. In some sense this is an expected result, since the restriction of $\Delta_{\gamma}$ to this subspace is a uniformly elliptic operator for which all the standard tools are available. Globally, the Grushin operator lacks a strong maximum principle for weak solutions and $W^{2,p}$-estimates which are used in order to derive additional regularity for weak solutions. We refer to Remark 6.8 for more details.
@@ 3 A functional setting for the Grushin operator
is the Sobolev critical exponent associated to $N_{\gamma}=m+(1+\gamma)\ell$, the homogeneous dimension associated to the decomposition $N=m+\ell$. Since the Sobolev-Gagliardo-Niremberg inequality proved in [29, Lemma 5.1]
@@
Let $\{u_{j}\}$ be a bounded sequence in $D_{\gamma}(\mathbb{R}^{N})$. We can suppose that $u_{j}\rightharpoonup 0$. For each $R>0$, we have the continuous embedding $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow\mathscr{H}_{\gamma}(B_{R}(0))$ [^3] For an open bounded set $\Omega\subset\mathbb{R}^{N}$, we denote by $\mathscr{H}_{\gamma}(\Omega)$ the completion of $C^{1}_{0}(\Omega)$ with respect to the norm $\lVert u\rVert_{\gamma}=\left(\int_{\Omega}\lvert\nabla_{\gamma}u\rvert^{2}\,{\mathrm{d}z}\right)^{1/2}.$ Since the embedding $\mathscr{H}_{\gamma}(B_{R}(0))\hookrightarrow L^{q}(B_{R}(0))$ is compact, see [26, Proposition 3.2], it follows that $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow L^{q}(B_{R}(0))$ is a compact embedding as well. Hence,
@@
for some subsequence. As $D_{\gamma}(\mathbb{R}^{N})\hookrightarrow L^{2^{*}_{\gamma}}(\mathbb{R}^{N})$ is a continuous embedding, we have that $\{\lvert u_{j}\rvert^{q}\}$ is a bounded sequence in $L^{2^{*}_{\gamma}/q}(\mathbb{R}^{N})$. By the Brezis-Lieb Lemma, [10, Theorem 1], up to a subsequence if necessary
@@ 4 Existence of solutions for (SP)
Now we are in a position to apply the Mountain Pass Theorem [3, Theorem 2.1], which ensures the existence and uniform boundedness of solutions to the (SP) problem.
@@
Consider $a_{1}$, $\delta$ and $v$ as given in Lemma 4.1. For $a\in(0,a_{1})$, using Lemma 4.1 and 4.2, the hypotheses of the mountain pass theorem [3, Theorem 2.1] are verified. So there exists a non-trivial critical point $u_{a}\in D_{\gamma}(\mathbb{R}^{N})$ of $I_{a}$ satisfying
@@ 5 Properties of the solutions
Our next result establishes that the solution $u_{a}$ belongs to $L^{\infty}(\mathbb{R}^{N})$ if the parameter $a$ is sufficiently small. Using the iterative argument of [32, Lemma B.3] we prove that the solution $u_{a}\in L^{q}(\mathbb{R}^{N})$, for every $q\in[2^{*}_{\gamma},\infty)$, provided $a$ is small enough. Moreover the argument in [30, Theorem 3.1] can be applied to establish the boundedness of $u_{a}$.
@@
Step 3. The continuity of the solutions $u_{j}$ comes from the non-homogeneous Harnack inequality for $X$-elliptic operators established in [22, Theorem 5.5] combined with some standard arguments, see for example [23, Corollary 4.18] for further details. ∎
@@ 6 Positive solutions
Let $\Sigma=\{0\}\times\mathbb{R}^{\ell}$ be the degenerate set for the Grushin operator. If $\Omega$ is any bounded domain contained in $\mathbb{R}^{N}\setminus\Sigma$, then $\Delta_{\gamma}$ is a uniformly elliptic operator on $\Omega$. Hence $u\in C^{2}(\Omega)$ and either $u\equiv 0$ on $\Omega$ or $u>0$ in $\Omega$. In the former case, it follows from [26, Proposition 2.5] that $u\equiv 0$ on the connected component of $\mathbb{R}^{N}\setminus\Sigma$ which contains $\Omega$. Therefore $u\equiv 0$ on $\mathbb{R}^{N}\setminus\Sigma$.
Now, it is clear that $\Sigma$ has empty interior, and it follows from [25, Theorem 9, Chapter 1] that $\mathbb{R}^{N}\setminus\Sigma$ is dense in $\mathbb{R}^{N}$. Recalling that $u$ is a continuous function, we conclude that $u\equiv 0$ on $\Sigma$. To this stage the function $u$ might be zero in (the interior of) some connected component of $\mathbb{R}^{N}\setminus\Sigma$ and strictly positive in (the interior of) some other component. However this is again impossible, since [26, Proposition 2.5] applies again on any ball which meets two such components across $\Sigma$. Hence either $u\equiv 0$ on $\mathbb{R}^{N}$, or $u>0$ on $\mathbb{R}^{N}$. The former is not admissible due to the Lemma 5.3. ∎
@@
When $\gamma$ is an even integer, a very elegant proof of the positivity of $u$ can be devised. We follow an idea of [6]. When $\gamma$ is an integer, it is known that $-\Delta_{\gamma}$ satisfies the Hörmander condition. Recalling that $g>0$ and $f_{0}\geq 0$, we see that $\int_{\mathbb{R}^{N}}\nabla_{\gamma}u\nabla_{\gamma}\varphi\,{\mathrm{d}z}\geq 0$ for every $\varphi\in C_{0}^{\infty}(\mathbb{R}^{N})$ such that $\varphi\geq 0$. Hence $\int_{\mathbb{R}^{N}}u(-\Delta_{\gamma}\varphi)\,{\mathrm{d}z}\geq 0$ for every such $\varphi$. Recalling that $u$ is a continuous function, it follows that $u$ is a viscosity subsolution of the operator $-\Delta_{\gamma}$. The strong maximum principle proved in [4, Corollary 1.4] yields that either $u\equiv 0$ or $u>0$ in $\mathbb{R}^{N}$. The first option, again, is not admissible due to the Lemma 5.3.
Now we are going to show that $u_{j}\to u$ in $L^{\infty}(\{0\}\times\mathbb{R}^{\ell})$. We essentially follow [1, 7], although a few important differences will arise.
@@
is a fundamental solution for the Grushin operator with singularity at $z=0$, see [5] or [16, Appendix B].
@@
$C^{-1}=(N_{\gamma}-2)\int\limits_{\{z\mid d(z)=1\}}\frac{\lvert x\rvert^{2\gamma}}{\left(\lvert x\rvert^{2\left(1+2\gamma\right)}+\left(1+\gamma\right)^{2}\lvert y\rvert^{2}\right)^{1/2}}\,\mathrm{d}S,$
@@
Let us consider a spherical shell $\Omega=B_{R_{2}}\setminus\overline{B_{R_{1}}}$, with $0\leq R_{1}<R_{2}\leq+\infty$. If $u$ is a radial function, by [15, Eq. (2.4)] we get the formula
@@
$\lVert u_{j}-u\rVert_{L^{\infty}(\{0\}\times\mathbb{R}^{\ell})}\to 0$ as $j\to\infty$.
@@
By [18, Sect. 3.1] and recalling that $u$ is a continuous function, we have the representation formula
$u(z)=C\int_{\mathbb{R}^{N}}\frac{g(\xi)f_{0}(u(\xi))}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi\quad\hbox{for a.e. $z\in\{0\}\times\mathbb{R}^{\ell}$}.$ (33)
@@
$u_{j}(z)=C\int_{\mathbb{R}^{N}}\frac{g(\xi)f_{j}(u_{j}(\xi))}{d(z-\xi)^{N_{\gamma}-2}}\,\mathrm{d}\xi\quad\hbox{for a.e. $z\in\{0\}\times\mathbb{R}^{\ell}$}.$ (34)
Using (33) and (34) we estimate $\lvert u-u_{j}\rvert$ as follows: for every $z\in\{0\}\times\mathbb{R}^{\ell}$,
@@
Therefore, (35) yields $u_{j}\to u$ in $L^{\infty}(\{0\}\times\mathbb{R}^{\ell})$ as $j\to\infty$. ∎
@@
$u_{j}\geq 0$ a.e. in $\{0\}\times\mathbb{R}^{\ell}$ for $j$ large enough.
@@
If (f3) and (g1) are satisfied then $u_{j}>0$ a.e. in $\{0\}\times\mathbb{R}^{\ell}$.
@@
for some $M>0$. For $z\in\{0\}\times\mathbb{R}^{\ell}\setminus\{(0,0)\}$, using (33) and (34), we get
@@
for any $R>0$. For any $R>0$ there exists $z\in\{0\}\times\mathbb{R}^{\ell}$ such that $d(z)>2R+1$. Hence
@@
Therefore, by (37) there exists $j_{2}\in\mathbb{N}$ and $R>>1$ such that for $j\geq j_{2},\,u_{j}>0$ a.e. on ($(\{0\}\times\mathbb{R}^{\ell})\setminus B_{R}$. Moreover, since $u$ is continuous, there exists $\eta>0$ such that $u>\eta$ on $\overline{B_{R}}$. Therefore, from Proposition 6.5, there exists $j_{3}\in\mathbb{N}$ such that for $j\geq j_{3},\,u_{j}>0$ a.e. on $\overline{B_{R}}\cap(\{0\}\times\mathbb{R}^{\ell})$. Thus, by choosing $j_{4}=\max\left\{j_{2},j_{3}\right\}$, we see that for $j\geq j_{4},\,u_{j}>0$ a.e. on $\{0\}\times\mathbb{R}^{\ell}$. This completes the proof. ∎
Remark 6.8.
The validity of the previous argument relies heavily on the representation formula (33), which is typically derived from the existence of a Green function associated to the differential operator. For linear differential operators with constant coefficients, the existence of a fundamental solution allows one to construct (continuous) solutions via convolution. by exploiting the invariance of the operator under translations. The Grushin operator $\Delta_{\gamma}$ is a differential operator with variable coefficients, and in particular it is invariant under translations only in the subspace $\{0\}\times\mathbb{R}^{\ell}$ of $\mathbb{R}^{N}$. So, even if a fundamental solution with singularity at a fixed point is available, it is not trivial to use it in order to derive an explicit Green function. As we have seen, it was proved in [18] that (33) holds true at points of the form $(0,y)$ at which the weak solution $u$ is continuous. It is interesting to observe that the authors of [17] proved the equivalence of a representation formula like (33) and of a Liouville-type result for the operator $\Delta_{\gamma}$, but only for $C^{2}$ solutions. However a full regularity theory for $\Delta_{\gamma}$ is not available, this approach remains unapplicable.
Furthermore, the existence of an approximate Green function for $\Delta_{\gamma}$ was proved in [29] and exploited in [2]. Unfortunately the known properties of it are too weak for our purposes, and would lead to replace (g1) with a completely implicit assumption.